### v2

# Who Gets Cited Most? Benchmarking Long-Context Reasoning on Scientific Articles

Abstract
We introduce SciTrek, a novel question-answering benchmark designed to evaluate long-context reasoning capabilities of large language models (LLMs) using scientific articles. Current long-context benchmarks often focus on simple information retrieval tasks, or employ artificial contexts. SciTrek addresses these limitations by creating benchmark questions that require information aggregation and synthesis across multiple full-text scientific articles. The questions and their ground-truth answers are automatically generated by formulating them as SQL queries over a database constructed from article metadata (i.e., titles, authors, and references). These SQL queries provide explicit, verifiable reasoning processes that enable fine-grained error analysis on model answers, and the data construction scales to contexts of up to 1M tokens with minimal supervision. Experiments on open-weight and proprietary LLMs show that SciTrek poses significant challenges as the context length increases, with supervised fine-tuning and reinforcement learning offering only limited gains. Our analysis reveals systematic shortcomings of frontier LLMs’ ability to effectively perform numerical operations and accurately locate information in long contexts. [^1] Our dataset and code are available from https://github.com/oaimli/SciTrek
Question SQL Query Answer
What is the highest number of authors that any single article has? SELECT MAX(author_count) FROM articles 10
What is the word count of the titles of articles, sorted by the number of authors in ascending order? SELECT title_word_count FROM articles ORDER BY author_count ASC 9, 17, 5, 6, 9, 12
How many references do articles with exactly two authors have? SELECT reference_count FROM articles WHERE author_count = 2 16
What is the total number of words in the titles of all articles that have exactly 60 references? SELECT SUM(title_word_count) FROM articles WHERE reference_count == 60 13
What are the names of the authors who are either the first or second author of an article, listed in descending order of their position? SELECT author_name FROM article_author WHERE author_position < 2 ORDER BY author_position DESC N. Shazeer, A. Vaswani
How many articles are cited by other articles but do not cite any other articles? SELECT COUNT(*) FROM articles WHERE article_id NOT IN (SELECT article_id_citing FROM citing_cited) AND article_id IN (SELECT article_id_cited FROM citing_cited) 2
Table 1: The SciTrek benchmark: example questions with corresponding SQL queries and answers.
1 Introduction
Large language models (LLMs) show promise in accelerating scientific progress by assisting researchers in reviewing and synthesizing the growing body of literature (Pearson, 2024; Agarwal et al., 2025) and generating novel research ideas (Si et al., 2025; Baek et al., 2025). So much so that industry products such as Deep Research [^2] https://openai.com/index/introducing-deep-research/, Elicit [^3] https://elicit.com/, and Scite [^4] https://scite.ai/ have emerged as dedicated tools to aid users with complex, multi-step research tasks. Scientific workflows often require processing large inputs: full scientific articles, collections of references, or structured datasets. A model capable of processing 100K+ tokens could simultaneously analyze multiple articles, track long chains of reasoning, or connect experimental results to prior work.
While long-context language models (LCLMs) are highly relevant to scientific tasks, existing benchmarks for evaluating their capabilities predominantly focus on non-scientific texts (Kuratov et al., 2024; Maekawa et al., 2025; Bai et al., 2025; Yen et al., 2025). They mostly address simple information retrieval tasks (e.g., Needle-In-A-Haystack; Kamradt 2023; Hsieh et al. 2024) rather than information integration, i.e., reasoning across multiple documents to synthesize information and answer complex questions. Some question-answering benchmarks (Asai et al., 2024; An et al., 2024; Bai et al., 2025) require information integration but do not provide a typology of the reasoning steps involved, making it difficult to determine which capabilities are being assessed and why a model fails.
Another issue relates to scalability and extensibility. When substantial human effort is required to generate benchmark questions and their answers (e.g., Asai et al. 2024), scaling to longer contexts or different types of questions becomes prohibitive. To better emulate scientific research practice, it is also important to evaluate a range of skills in natural contexts, such as a model’s ability to aggregate or filter information, and combinations thereof (Wang et al., 2024c). In contrast, a few benchmarks rely on artificial contexts, created by adding noise  (Kuratov et al., 2024) to short contexts (e.g., HotpotQA; Yang et al. 2018) or by synthesizing the input with LLMs (Maekawa et al., 2025).
In this paper, we create a question-answering benchmark based on scientific articles that aims to alleviate these issues. We assume a scenario where a researcher, possibly during the literature review process, seeks answers to specific questions when the related scientific articles are provided. As an example, consider the first question in Table 1: “What is the highest number of authors that any single article has?”. Although seemingly simple, it requires aggregation of information across multiple articles. To answer correctly, a hypothetical model would need to count the authors of each article in the collection and select the one with the largest number. We generate questions requiring explicit reasoning by focusing on basic numerical operations such as counting, aggregating, and sorting, applied to core elements of scientific articles: titles, authors, and references. Since these operations correspond to standard database functions, we can automatically generate arbitrarily complex SQL queries over any collection of scientific articles and obtain ground-truth answers without manual annotation by constructing database tables for titles, authors, and references. We exploit recent advances in LLM capabilities for SQL understanding and generation (Yu et al., 2018; Li et al., 2023; Hui et al., 2024) to automatically convert SQL queries into natural-language questions.
Our benchmark, SciTrek, contains a test set of 2,121 question-answer pairs on scientific articles and a training set of 19,543 question-answer pairs, with contexts of varying lengths (i.e., 64K, 128K, 512K, and 1M). [^5] In the paper, K refers to 1,024 and M to 1,024K. Table 1 shows examples of questions, their corresponding SQL queries, and answers. While these questions are relatively superficial and do not engage deeply with the domain-specific content, they require models to retrieve and synthesize facts that are sparsely distributed across article collections and span long contexts. If models struggle with simple comparison, filtering, sorting, and aggregation operations, they might be impossible to succeed on more complex analytical tasks (Wolfson et al., 2020, 2025). We use SciTrek to evaluate the reasoning capabilities of LCLMs with context windows exceeding 128K tokens. Experimental results reveal that SciTrek poses significant challenges for these LCLMs. In general, models achieve poor performance, even when enhanced with supervised fine-tuning (SFT) and reinforcement learning (RL). This paper makes the following contributions:
  • We propose a methodology for constructing a long-context question-answering benchmark over natural texts, where explicit reasoning processes are approximated via SQL operations involving information aggregation. The methodology requires minimal human intervention and generalizes to any domain with entities that can be structured as database tables.
  • Extensive experiments reveal that frontier open-weight and proprietary LLMs struggle with this task, especially as input length increases. While SFT and RL improve over baseline models, performance remains limited. All models perform poorly on questions involving references and citation relationships among articles.
  • By design, SciTrek enables fine-grained analysis of model behavior, offering insight into where and why models fail. Evaluation of open-weight models reveals systematic shortcomings in counting bibliographic elements and performing basic numeric operations, with error rates increasing as question length grows; models also frequently misinterpret compound conditions and struggle with logical constructs involving negation (see last question in Table 1).
Dataset Natural Context #SciQs Reasoning Scalable Len     Example Question
NeedleBench 128K What legendary item is hidden on Emerald Island?
Ada-LEval 128K What is the correct order of the segments?
BABILong 10M Where is Mary?
HELMET 128K Who set the fire in one tree hill?
LIFBENCH 128K Retrieve the entry at position 8th in the list.
RULER 128K Find all variables that are assigned the value 12345.
OpenScholar 208 3K Compile a list of reviews […], and identify the most promising […]
LongBench v2 50 128K How long have I been living in my current apartment in Shinjuku?
LongMemEval 2M How many bikes do I currently own?
L-Eval 200K How do I know when I should apply for Medicare?
HoloBench 64K What are the names of wines and their corresponding grape types?
MathHay 128K What is the total number of points scored by LeBron […] combined?
Loong 53 250K Which company has the highest non-current assets?
SciTrek 1,716 1M How many articles are cited by other articles but do not cite any other articles?
Table 2: Representative benchmarks for evaluating LCLMs. Natural Context: does the benchmark have natural input contexts; #SciQs: the number of unique benchmark questions or instructions with ground-truth answers on scientific articles (—: benchmarks that are not using any scientific articles); Reasoning: does the benchmark provide detailed reasoning skills required to answer each question; Scalable: can the benchmark be extended with minimal effort (e.g., to longer contexts or larger datasets); Len: the maximum input length that the model supports in terms of tokens; : length based on texts retrieved rather than full input (which is not available).
2 Related Work
Long-context Language Modeling
A growing body of work explores the challenges LLMs encounter when processing inputs that exceed their context lengths. Rotary Position Embeddings (RoPE; Su et al. 2024) have emerged as a widely used approach that replaces absolute positional embeddings with rotational transformations of token embeddings, improving generalization to unseen sequence lengths. Building on RoPE, several extensions introduce rescaling to further enhance performance on substantially longer inputs without the need for retraining (Peng et al., 2024).
Other architectural advances focus on reducing the quadratic cost of the Transformer attention mechanism (Sun et al., 2025). Sparse attention methods, such as LongFormer (Beltagy et al., 2020), compute only selected portions of the full attention matrix, thereby reducing computational cost and accelerating both prefilling and inference (Jiang et al., 2024; Fu et al., 2024). Many contemporary models interleave sparse and full attention layers to balance efficiency and performance (Dubey et al., 2024; Llama Team, 2025; Yang et al., 2025; Kamath et al., 2025). Another popular strategy for reducing computation in long-context settings is to limit the number of active model parameters. Group Query Attention (GQA; Ainslie et al. 2023) enables multiple attention heads to share key–value projections, without significant performance loss. The open-weight LLMs used in our experiments combine these techniques, complemented by pretraining on long-context tasks (Dubey et al., 2024; Kamath et al., 2025; Yang et al., 2025; Llama Team, 2025).
Long-context Datasets
The growing interest in LCLMs has driven the development of numerous benchmarks designed to evaluate their performance. Initial benchmarks primarily assessed whether models can retrieve information from their context, typically through Needle-In-A-Haystack (NIAH) tasks, such as NeedleBench (Li et al., 2024). Building on these approaches, subsequent benchmarks either added distractor information to short-context tasks or scattered relevant information across synthetic contexts, such as Ada-LEval (Wang et al., 2024a), BABILong (Kuratov et al., 2024), HELMET (Yen et al., 2025), LIFBENCH (Wu et al., 2025b), LongBench (Bai et al., 2024), and RULER (Hsieh et al., 2024).
In addition to retrieval, some benchmarks emphasize complex reasoning over long contexts. For example, OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) contain questions that require extensive use of multiple documents provided as context, while DocFinQA (Reddy et al., 2024) and MedOdyssey (Fan et al., 2025) focus on financial and medical reasoning, respectively. Although not explicitly technical, LongMemEval (Wu et al., 2025a) and L-Eval (An et al., 2024) also require reasoning over natural, long-context information. However, evaluating model answers against these benchmarks remains challenging, as the reasoning steps required to arrive at correct responses are often opaque, making it difficult to assess failure modes. In contrast, HoloBench (Maekawa et al., 2025) and MathHay (Wang et al., 2024b) better delineate between the types of reasoning required for each question, but rely on unnatural contexts. CURIE (Cui et al., 2025), OpenScholar, and LongBench v2 evaluate reasoning over scientific articles, but their reliance on expert-curated questions and answers poses challenges for scalability to longer contexts or larger datasets.
In designing SciTrek, we aim to retain many of the useful qualities of previous benchmarks (e.g., structured reasoning, scalability) while still proposing a fairly natural and easy task for humans to perform. We compare SciTrek with existing benchmarks in Table 2. Unlike Loong (Wang et al., 2024c), which tests broad long-context comprehension, SciTrek uniquely targets scientific synthesis and structured reasoning at extreme context lengths. Although we focus on scientific articles, our methodology could extend to other domains with explicit entities and relations. Thanks to the SQL backbone, we are able to construct questions testing various model skills (e.g.,  aggregation, filtering), and how these manifest themselves through explicit reasoning processes.
3 The SciTrek Benchmark
Figure 1: Overview of SciTrek construction process: we gather article collections of varying scales; we then obtain SQL queries and their answers based on databases (which store collection-specific metadata); and finally we convert SQL queries to natural language questions. The final dataset consists of input contexts (full-text scientific articles), natural language questions, and their answers, all highlighted in blue.
Figure 1 illustrates how SciTrek was curated: we first gather scientific article collections as contexts corresponding to different lengths (i.e., 64K, 128K, 512K, and 1M); we then create databases representing the article metadata and SQL queries with answers from database execution; finally, we convert the SQL queries into natural language questions. We explain these stages in detail below.
3.1 Gathering Scientific Article Collections
SciTrek is constructed using scientific articles from Semantic Scholar. [^6] https://www.semanticscholar.org/ To cover diverse topics of articles, we obtain an initial set of seed articles from eight subjects: Computer Science (CS), Economics, Electronic Engineering (EE), Math, Physics, Biology, Finance, and Statistics. [^7] Following the typology of subjects from https://arxiv.org/. For each subject, we select two seed articles with more than 100 citations since 2020. For each article, we retrieve related articles from Semantic Scholar. [^8] Semantic Scholar provides APIs to retrieve scientific articles with metadata including titles, authors, a reference list with articles that the current article cites, and a citation list with articles that cite the current article. Throughout this paper, a reference is a bibliographic entry listed in the reference section of a scientific article, whereas a citation is an in-text acknowledgement of another’s work within the article. To ensure broad coverage, we include two-hop related articles identified via Semantic Scholar’s citation graph based on the reference and citation list. For each seed article, we randomly sample ten first-hop related articles, and for each of these, we further sample five second-hop related articles to form an article cluster. Since full texts are required, we filter out articles without PDFs. This process yields 16 article clusters comprising 662 scientific articles with PDFs across eight subjects (see step I in Figure 1). Finally, we convert the collected PDFs into markdown texts using Marker. [^9] https://github.com/datalab-to/marker
From these clusters, we create article collections of varying context lengths for our question-answering task. By concatenating the markdown texts, we generate collections with lengths of 64K, 128K, 512K, and 1M tokens. [^10] We can easily construct collections with arbitrary lengths exceeding 1M tokens. Each collection is initialized with a randomly selected article, and additional articles are added until it reaches a specified length level. We expand the collections using two strategies: (1) random sampling from the clusters, and (2) traversing the citation graph within each cluster (using both depth-first and breadth-first search) to construct collections that preserve citation relations among articles. Each collection contains at least four articles, and no two collections share more than half of their articles. From the 662 scientific articles across 16 clusters, we construct 2,612 article collections spanning all four length levels, including 2,027 generated through random sampling and 585 through graph traversal.
3.2 Creating Databases and SQL Queries
Once the article collections are assembled, we construct a database for each (see step II in Figure 1), focusing on a core subset of elements shared across all articles: titles, authors, and references. Based on these key elements, each database comprises three tables: articles, article-author, and citing-cited. The articles table contains metadata such as the title, reference count, and title word count. The article-author table captures author information, including names and their positions in the corresponding author list. The citing-cited table records citation relations among articles. A detailed description is provided in Appendix A. This information is obtained from Semantic Scholar or derived via simple preprocessing, such as splitting titles and counting words. Manual corrections were applied as needed to align with the full-text Markdown articles.
SQL Commands
Aggregating MAX, MIN, SUM, AVG, COUNT, DISTINCT
Sorting ORDER BY, ASC, DESC, GROUP BY
Filtering WHERE
SQL Operators
Comparison ==, >>, <<, >=>=, <=<=, <><>, LIKE
Arithmetic ++, -, *, /, %
Logical AND, NOT, OR, BETWEEN, IN
Table 3: Core commands and operators in SQL.
Skill Count     Example Query Template
Aggregating 20 SELECT MAX(author_count) FROM articles
Sorting 27 SELECT title_wstatisticsord_count FROM articles ORDER BY author_count ASC
Filtering 107 SELECT author_name FROM article_author WHERE author_position == {author-position}
Filtering+ Aggregating 107 SELECT SUM(title_word_count) FROM articles WHERE reference_count == {reference-count}
Filtering+ Sorting 106 SELECT author_count FROM articles WHERE title_word_count % 2 == 1 ORDER BY title_word_count DESC
Relational Filtering 20 SELECT COUNT(*) FROM articles WHERE article_id NOT IN (SELECT article_id_citing FROM citing_cited) AND article_id IN (SELECT article_id_cited FROM citing_cited)
Table 4: SQL templates representing different information processing skills. {author-position}, {reference-count} are placeholders.
The core SQL [^11] Structured Query Language (SQL) is a standardized language for managing and querying relational databases. commands summarized in Table 3 form the foundation for building SQL queries, often combined with operators (comparison, arithmetic, logical) to create more complex queries (since WHERE is used to filter data, it always works in conjunction with SQL operators). Aside from the SQL commands in Table 3 (Aggregating, Sorting, and Filtering), we define composite commands based on their combinations (i.e., Filtering+Aggregating and Filtering+Sorting). Using these, we manually create SQL query templates targeting different topics related to key elements of scientific articles, including Author Count, Author List, Reference Count, Title List, and Title Word Count. This process is illustrated in Figure 1 (right panel) and example templates are shown in Table 4; Note that some templates include placeholders to be instantiated with specific values. Finally, to capture authorship and citation relations, we introduce Relational Filtering and the topics of Author Relation and Citation Relation, which specifically target authorship and citation relations.
We collectively refer to the SQL commands in Table 4 as information processing skills, since they test different information processing capabilities. We have various templates per skill, designed to be applicable across article collections. For each collection, we randomly select 10 templates and instantiate them with collection-specific values for all placeholders. We then execute these queries against the corresponding database to generate ground-truth answers.
3.3 Converting SQL Queries to Natural Language
We use Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) to convert SQL queries into natural language questions. To ensure queries and questions are meaning-preserving, we validate each generated question by converting it back to SQL and verifying that both queries produce identical results when executed against the collection database. For each SQL-collection pair, we repeat this process up to 10 times; if no valid question is obtained, we discard the query for that collection. Using the prompts in Appendix E, we successfully generate natural language questions for 82.9% of SQL-collection pairs. Following this, we obtain 2,121 test questions for evaluating models against the four context lengths defined in Section 3.1, and 19,543 instances for training (see Section 4.2). Table 5 presents descriptive statistics for the SciTrek test set. For each context length, our test partition covers all information processing skills and question topics described in Section 3.2. The distribution of SQL commands and operators is provided in Appendix A, Table 11.
3.4 Data Quality Validation
Length Instances Articles/C Words/Q Words/A
64K 112 4.2 14.6 6.2
128K 728 5.7 16.9 7.9
512K 667 22.7 18.1 30.9
1M 614 46.3 18.6 63.6
Table 5: Descriptive statistics for SciTrek (test set). Length: the length level of input tokens; Instances: number of question-answer pairs with their contexts; Articles/C: average number of scientific articles per context; Words/Q: average number of words per question; Words/A: average number of words per answer.
To validate SciTrek’s quality, we crowdsourced human annotations through Prolific. [^12] https://www.prolific.com Annotators (English native speakers from the US or the UK) were asked to answer natural language questions from a random sample of 120 instances, with 20 instances representing each of the six skills listed in Table 4. Crowdworkers answered questions using database tables rather than full-text articles, as documents spanning 1M tokens are impractical for humans to review.
Each instance was annotated by three participants who were given the same instructions as those used for model testing with database tables as context (Figure 7, Appendix E). In each annotation session, crowdworkers were given three questions with different contexts. To ensure annotation quality, each session included a quality control question. On average, annotators spent about 5.5 minutes per session and were compensated above the UK living wage, at £12 per hour.
We measured inter-annotator agreement and the alignment between database-derived answers (from SQL query execution) and human responses using exact match. Annotators were considered in agreement if two or more provided the same answer. Similarly, database answers were considered aligned with human responses if they matched the responses of two or more annotators. Table 6 reports average agreement and alignment across 120 instances, showing strong inter-annotator agreement and high consistency between database-derived and human answers.
Manual inspection of all 120 questions given to annotators reveals that 3 of the 19 Relational Filtering questions were ambiguous, explaining the gap between alignment and agreement for this category in Table 6. While Qwen2.5-Coder-32B-Instruct generally converts SQL queries to natural language accurately, the complexity of Relational Filtering queries can introduce ambiguity.
Skill Agree (%) Align (%)
Aggregating 85.7 85.7
Sorting 85.0 80.0
Filtering 85.7 85.7
Filtering+Aggregating 95.0 85.0
Filtering+Sorting 89.5 89.5
Relational Filtering 89.5 73.7
All 88.3 83.3
Table 6: Inter-annotator agreement measured on a sample of 120 answers. Answers obtained from executing SQL queries are considered aligned if they match those provided by two or more annotators.
Full-text Articles Database Tables
Model Context Size F-64 F-128 F-512 F-1024 D-64 D-128 D-512 D-1024
Qwen2.5-7B-Instruct-1M 1M 4.5 2.8 0.3 0.0 20.5 14.3 5.0 2.0
Qwen2.5-14B-Instruct-1M 1M 8.3 6.5 1.6 0.1 33.3 27.2 11.0 5.9
Qwen3-4B-Instruct-2507 256K 2.1 7.2 25.9 16.9 6.9 2.8
Qwen3-4B-Thinking-2507 256K 41.1 29.3 90.2 83.1 71.8 52.3
Qwen3-30B-A3B-Instruct-2507 256K 5.4 3.2 29.8 21.2 6.0 4.2
Qwen3-30B-A3B-Thinking-2507 256K 53.3 42.0 92.3 86.2 73.5 61.1
Gemma-3-27B-IT 128K 6.2 3.4 31.8 25.0 11.3 6.1
Llama-4-Scout-17Bx16E-Instruct 10M 5.4 2.8 1.3 1.1 28.6 19.0 7.4 4.0
Llama-3.3-70B-Instruct 128K 8.3 3.5 47.0 36.2 14.6 8.1
DeepSeek-R1-Distill-Llama-70B 128K 22.0 6.0 83.3 74.2 56.8 42.2
Gemini 2.5 Pro 1M 41.7 26.0 $\star$ $\star$ 91.7 83.5 55.4 31.5
GPT-4.1 1M 21.1 11.7 3.9 2.5 69.3 53.8 24.8 16.1
o4-mini 195K 61.0 46.5 95.2 87.8 79.4 72.6
Table 7: Evaluation of open-weight (top) and proprietary (bottom) models on SciTrek using exact match (%). Results are reported for two settings: using the full-text scientific articles as context and using the corresponding database tables. F-* refers to test data showing full-text article clusters with different maximum lengths, 64K/128K/512K/1M. D-* refers to test data showing the database tables corresponding to F-64/128/512/1024 as input. Note that for any given article cluster, the database representation is significantly shorter (mean length $\sim$}2K tokens); therefore we are able to pass database representations for all lengths to all models. Context Size denotes the maximum context length supported by each model. — indicates the model cannot handle the given context size; $\star$ indicates that the model was not evaluated due to prohibitive computational cost.
4 Results
In our experiments, we evaluate both proprietary and open-weight models from various families that support contexts exceeding 128K and have shown strong performance on language understanding and mathematical reasoning benchmarks, e.g., MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b). For each family, we select the largest model we can feasibly run with our resources. [^13] Experiments were conducted on 4 NVIDIA HGX H200s. For example, we use Llama-4-Scout-17Bx16E-Instruct instead of Llama-4-Maverick-17B-128E-Instruct, and the distilled variant of DeepSeek-R1. Table 7 provides an overview of the models we consider, which vary in parameter scale and supported context length (most were released within the past six months). For detailed model descriptions and settings, refer to Appendix B. We evaluate models in two context settings: (1) using the full-text scientific articles within a given collection as context and (2) using only the corresponding database tables. All models in our study are instruction-tuned, and we employ a uniform set of prompts across them, as detailed in Appendix E. We assess performance using average exact match and F1, as the expected outputs are factual items with minimal variation, such as specific numbers, author names, or article titles. For higher confidence, reported metrics are averages across three runs.
4.1 Zero-shot Prompting
Our zero-shot results are summarized in Table 7. Overall, we observe that SciTrek is challenging especially when using the full-text articles as contexts. In this setting, performance drops for all models as the input gets longer. This trend also manifests itself when using database tables as context. Perhaps unsurprisingly, proprietary models significantly outperform open-weight ones. We observe similar tendencies when using F1 as the evaluation metric (see Appendix C).
Figure 2 presents a fine-grained analysis for the best six models across question topics (e.g., author count, reference count, citation relations), skills (e.g., aggregating, sorting), and subjects (e.g., Economics, Biology) when using full-text articles as context. Model performance shows little variation by subject. However, most models struggle more with sorting tasks, while performing better on aggregation. Performance is lowest on citation-related questions (i.e., Citation Relation, Reference Count), and somewhat higher on author- and title-related questions.
Figure 2: Fine-grained analysis for best performing models in terms of exact match (EM). We examine how model performance varies across question topics (top), information processing skills (middle) and subjects (bottom) when using full-text scientific articles as context with an input length of 128K.
4.2 Long-context Post-training
Leveraging the data generation methodology described in Section 1, we curate a substantial training dataset (19,543 instances) across four context lengths including 64K, 128K, 512K, and 1M. We use this data to assess whether more training could improve the performance of open-weight models on full-text articles. We experiment with two well-established techniques: supervised fine-tuning (SFT) and reinforcement learning (RL). Because of the high computational cost, for this suite of experiments, we only report results with Qwen2.5-7B-Instruct-1M on data with maximum context of 128K tokens (7,703 training instances).
Moreover, we examine whether the post-trained model can generalize along the dimensions of input length, question topics, and information processing skills. We compare in-distribution performance where the model is trained and tested in similar conditions (e.g., on lengths of 64K and 128K) to out-of distribution where the model is tested on an unseen dimension. Specifically, we train the model on data corresponding to 64K and 128K context lengths and evaluate it both in-distribution (64K, 128K) and out-of-distribution (512K, 1M). For question topic, training uses author- and title-related questions at 64K and 128K, with evaluation covering the same topics in-distribution and citation-related questions out-of-distribution. For skills, training is performed on non-relational questions (Aggregating, Sorting, Filtering, Filtering+Sorting, Filtering+Aggregating), with evaluation including these in-distribution skills and the out-of-distribution Relational Filtering skill.
For SFT, we train the model for 500 steps with a batch size of 32, a learning rate of $2\times 10^{-6}$, and a warm-up rate of 0.05. For RL, we use GRPO (Shao et al., 2024) with a mixed reward of EM and F1 to encourage the model to produce both reasoning traces and answers. GRPO was chosen due to its success in similar verifiable and long-context tasks (Shao et al., 2024; Mroueh, 2025; Gurung and Lapata, 2025; Zheng et al., 2025). We found a simple sum of EM and F1 to improve accuracy over time while still providing a useful training signal from difficult questions. The prompt for generating reasoning traces is provided in Appendix E. Since GRPO optimization is time-consuming, we restrict training to a single epoch which takes around 5 days.
Length Topic Skills
Models ID OOD ID OOD ID OOD
Qwen2.5 (ZS) 3.1 0.2 3.9 1.5 1.4 5.8
Qwen2.5 (SFT) 16.3 2.3 20.9 10.0 10.8 19.2
Qwen2.5 (GRPO) 22.5 2.0 30.6 7.5 20.0 26.8
GPT-4.1 (ZS) 13.0 3.4 17.3 5.0 8.0 21.5
Gemini 2.5 Pro (ZS) 28.1 36.0 13.5 24.9 33.7
o4-mini (ZS) 48.4 62.5 22.3 53.7 39.3
Table 8: Post-training Qwen2.5-7B-Instruct-1M on SciTrek. Results with SFT and GRPO are compared with zero-shot prompting (ZS) of the same model and other proprietary models in terms of exact match (%). ID is a shorthand for in-distribution, and OOD for out-of-distribution. —: the model cannot handle the context size or has prohibitive computational cost.
The results in terms of exact match in Table 8 demonstrate that both SFT and GRPO slightly improve performance across dimensions, surpassing GPT-4.1 despite its much larger parameter size. However, Qwen2.5-7B-Instruct-1M is still not able to generalize to longer inputs, but does improve on out-of-distribution question topics and information processing skills.
5 Model Failure Analysis
Model #Articles Question Length SQL Length Ground-truth Answer Length
Qwen2.5-7B-1M -0.04 -0.15   0.11 -0.01
Qwen2.5-14B-1M -0.04 -0.16   0.07 -0.06
Qwen3-4B-Instruct-2507 -0.01 -0.16   0.05 -0.03
Qwen3-4B-Thinking-2507 -0.06 -0.14   -0.19 -0.01
Qwen3-30B-A3B-Instruct-2507 -0.02 -0.15   0.06   0.08
Qwen3-30B-A3B-Think-2507   -0.12 -0.08   -0.19 -0.02
Gemma-3-27B-IT -0.08 -0.14   0.02 -0.08
Llama-4-Scout   -0.10 -0.17 -0.03 -0.03
Llama-3.3-70B -0.06 -0.16 -0.03 -0.05
DeepSeek-R1-Distill-Llama-70B   -0.15 -0.16 -0.02 -0.08
Gemini 2.5 Pro   -0.16 -0.22 -0.03 -0.01
GPT-4.1 -0.07 -0.21   0.06   0.06
o4-mini   -0.14 -0.09   -0.25 -0.06
Qwen2.5-7B-Instruct-1M (SFT) -0.03 -0.32   0.09   -0.14
Qwen2.5-7B-Instruct-1M (GRPO) -0.07 -0.13 -0.02 -0.08
Table 9: Pearson correlation between various factors and model performance using exact match (with full-text articles as context, 128K token length). The length of questions, SQL queries, and ground-truth answers is computed by counting words separated by spaces. Bold values indicate the strongest correlation for each model, and denotes significant correlations with p-value <<0.05. Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block.
We conduct additional analysis to understand model failures on SciTrek when using full-text articles as context. Table 9 shows correlations between model performance and various factors: the number of input articles, the question length (LenQ), the length of the underlying SQL query (LenSQL), and the answer length (LenA). They indicate that model performance correlates primarily with the question length, while o4-mini shows correlations with the number of input articles and the length of the SQL query.
Inspection of model answers reveals several failure patterns: (1) weaker models tend to respond by simply outputting “NULL”, which suggests they rely more on the instruction to produce a fallback answer rather than genuinely understanding the given context; (2) models often fail to follow the specified output formats, e.g., returning author lists when author numbers are requested; (3) models sometimes provide incomplete answers when lists are requested for questions that require aggregation; and (4) models misinterpret compound conditions, particularly struggling with negation-based filtering (e.g., the last example in Table 1). More details of failure analysis are provided in Appendix D.
We also manually analyzed 200 reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates. [^14] Reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B are lengthy, averaging 638, 1,685, and 2,320 respectively. We will release these traces along with our code, but we omit them here due to space constraints. We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis. Specifically, we examined whether the reasoning is logically sound and whether individual steps obtain accurate information from specific contexts. We find that (1) the overall reasoning remains logically sound at 64K and 128K context lengths for all three models, although their traces occasionally include repeated steps or extraneous operations unrelated to the question; (2) despite not being trained for 512K, Qwen2.5-7B-1M (GRPO) still demonstrates coherent reasoning, but with reduced accuracy in specific steps compared to 64K and 128K; and (3) most reasoning steps that require counting are not accurate especially for references, which leads to incorrect answers from these models especially for Qwen2.5-7B-1M (GRPO). These findings indicate that GRPO-based reinforcement learning improves abstract reasoning without enhancing the model’s accuracy in fine-grained operations, such as counting references.
6 Conclusion
We introduce SciTrek, a benchmark for evaluating LLMs’ ability to perform multi-document information synthesis and structured reasoning over full-text scientific articles. By generating questions and ground-truth answers through a SQL backbone operating on article metadata, we provide a highly scalable framework with explicit reasoning processes that enables fine-grained error analysis.
Our extensive evaluation demonstrates that SciTrek poses a significant challenge to open-weight and proprietary LLMs, with only modest performance gains observed through supervised fine-tuning and reinforcement learning. Specifically, models struggle with compound logical conditions, fail at sorting tasks, and often produce incomplete or poorly formatted outputs. We believe the proposed methodology for creating SciTrek generalizes beyond titles, authors, and references to encompass broader metadata elements and domains, offering a tool to diagnose persistent shortcomings in LCLMs and improve their capability to support scientific workflows.
7 Limitations
While SciTrek benchmarks fundamental reasoning capabilities (including filtering, sorting, aggregating and their combinations) of LCLMs on scientific articles, it covers only a limited set of core elements of scientific articles, specifically titles, authors, and references. Future work will explore additional long-context model capabilities for understanding scientific articles, such as interpreting figures and performing domain-specific content reasoning.
Appendix A Details on SciTrek Construction
We construct a database with three tables for each article collection, using metadata from Semantic Scholar along with basic preprocessing. A description of these tables is provided in Table 10.
Our test dataset covers all SQL commands and operators that are listed in Table 3. The detailed distribution is shown in Table 11.
Database Column Data
Table Name Type Description
articles article_\_id String the unique identifier of the article
article_\_title String the title of the article
title_\_word_\_count Integer the number of words in the article’s title (using spaces to determine word boundaries)
author_\_count Integer the number of authors in the article
reference_\_count Integer the number of references that are cited in the article
article-author relation_\_id String the unique identifier of the article-author relations
article_\_id String the identifier of the associated article
author_\_name String the name of the author/s
author_\_position Integer the position of the author in the author list (starting from 0 for the first author)
citing-cited relation_\_id String the unique identifier of the citation relations between two articles
article_\_id_\_citing String the identifier of the article that cites another article
article_\_id_\_cited String the identifier of the article that is cited by another article
Table 10: Description of the database tables that we use to curate SQL queries and answers.
Command/ Command/
Operator Count % Operator Count %
SELECT 2,121 100.00 GROUP BY 175 8.25
WHERE 1,821 85.86 $\leq$ 175 8.25
== 1,032 48.66 % 164 7.73
IN 682 32.15 $\geq$ 154 7.26
OR 616 29.04 NOT 154 7.26
ORDER BY 591 27.86 $\neq$ 153 7.21
<< 476 22.44 AVG 135 6.36
>> 450 21.22 MIN 133 6.27
COUNT 370 17.44 BETWEEN 106 5.00
ASC 301 14.19 SUM 102 4.81
DESC 287 14.19 // 100 4.71
DISTINCT 280 13.20 LIKE 50 2.36
* 215 10.14 ++ 48 2.26
MAX 213 10.04 - 23 1.08
AND 195 9.19
Table 11: Distribution of SQL commands and operators in our test data. $\%$ refers to the prevalence of each command/operator in our test data’s SQL queries. These percentages do not need to add to 100%.
Appendix B Experimental Settings
Configuration details for the models we use are presented in Table 12. For all models, we use their default inference configuration for answer generation from their Huggingface repositories. We include both reasoning models and non-reasoning models. Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, DeepSeek-R1-Distill-Llama-70B, Gemini 2.5 Pro and o4-mini are reasoning models. In our experiments, o4-mini conducts reasoning with the default thinking effort medium, Gemini 2.5 Pro was given a thinking budget of 512 tokens, and Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B were prompted with the reasoning prompt in Figure 6.
Model #Parameters Open Weights Context Length Release Date
Qwen2.5-7B-Instruct-1M 7B 1,010,000 Jan 2025
Qwen2.5-14B-Instruct-1M 14B 1,010,000 Jan 2025
Qwen3-4B-Instruct-2507 4B 262,144 July 2025
Qwen3-4B-Thinking-2507 4B 262,144 July 2025
Qwen3-30B-A3B-Instruct-2507 3Bx10E 262,144 July 2025
Qwen3-30B-A3B-Thinking-2507 3Bx10E 262,144 July 2025
Gemma-3-27B-IT 27B 131,072 Mar 2025
Llama-4-Scout-17Bx16E-Instruct 17Bx16E 10,485,760 Apr 2025
Llama-3.3-70B-Instruct 70B 131,072 Dec 2024
DeepSeek-R1-Distill-Llama-70B 70B 131,072 Jan 2025
Gemini 2.5 Pro 1,048,576 Jun 2025
GPT-4.1 1,047,576 Apr 2025
o4-mini 200,000 Apr 2025
Table 12: Proprietary and open-weight LCLMs that are evaluated in our experiments. (Parameter numbers for proprietary models are not accessible. “17B×16E" indicates that the model uses a mixture-of-experts architecture with 16 experts, each containing 17 billion parameters. Context length is based on the maximum number of tokens these models can handle as input.)
Full-text Articles Database Tables
Model Context Size F-64 F-128 F-512 F-1024 D-64 D-128 D-512 D-1024
Qwen2.5-7B-Instruct-1M 1M 8.7 7.0 1.1 0.0 39.2 39.3 28.7 21.2
Qwen2.5-14B-Instruct-1M 1M 20.6 17.5 7.4 1.0 58.2 53.9 44.8 34.8
Qwen3-4B-Instruct-2507 256K 7.2 7.3 47.2 46.0 38.9 31.1
Qwen3-4B-Thinking-2507 256K 50.2 42.3 92.7 85.8 79.4 63.6
Qwen3-30B-A3B-Instruct-2507 256K 15.0 16.4 55.6 53.8 44.7 42.0
Qwen3-30B-A3B-Thinking-2507 256K 62.5 55.1 94.4 88.3 80.8 74.5
Gemma-3-27B-IT 128K 22.6 14.7 52.8 52.5 47.8 40.7
Llama-4-Scout-Instruct 10M 20.3 17.4 15.6 14.8 49.8 44.9 40.5 37.4
Llama-3.3-70B-Instruct 128K 25.3 15.1 65.8 63.5 53.7 49.5
DeepSeek-R1-Distill-Llama-70B 128K 29.0 9.0 85.5 77.6 68.0 62.3
Gemini 2.5 Pro 1M 58.1 48.8 $\star$ $\star$ 95.3 88.3 79.0 69.8
GPT-4.1 1M 36.0 29.7 22.3 19.6 82.8 73.3 61.8 56.3
o4-mini 195K 74.0 63.8 97.2 89.2 87.0 86.5
Table 13: Evaluation of open-weight (top) and proprietary (bottom) models on the SciTrek benchmark using F1 (%). Results are reported for two settings: using the full-text scientific articles as context and using the corresponding database tables. F-* refers to test data showing full-text article clusters with different maximum lengths, 64K/128K/512K/1M. D-* refers to test data showing the database tables corresponding to F-64/128/512/1024 as input. Note that for any given article cluster, the database representation is significantly shorter (mean length $\sim$}2K tokens); therefore we are able to pass database representations for all lengths to all models. Context Size denotes the maximum context length supported by each model. — indicates the model cannot handle the given context size; $\star$ indicates that the model was not evaluated due to prohibitive computational cost.
Appendix C Additional Results Using F1
Table 13 presents additional results using F1 as the evaluation metric. Models are tested against full-text scientific articles and database tables as context. The input length averages only 1,980 tokens when database tables of are used as context across different lengths of full-text article collections.
Appendix D Details on Model Failure Analysis
To gain deeper insight into model behaviour, we analyzed the failure patterns exhibited in model-generated answers on SciTrek. We randomly selected 60 instances in the input length of 128K that were hard for all zero-shot and supervised models (10 instances per skill in Table 4). We find that models tend to output “NULL" particularly on samples that require the skill of filtering. As shown in Table 14, most open-weight models tend to randomly output “NULL" for any sample. This indicates that these models rely more on the instruction to produce a fallback “NULL” answer rather than on genuine understanding of the provided context. [^15] Our instruction prompted models to generate “NULL” when they could not get the answer from the context. In addition, we find that models fail to follow the specified format for the answers, especially for sorting-related questions, as shown in Table 15. For example, models tend to generate author lists when sorted author numbers are requested and output lists when aggregates are requested. Interestingly, GPT-4.1 produces incorrectly formatted answers far more frequently than all other models. We further observe that some models tend to produce partial answers when lists are requested, especially for tasks that require aggregation. As shown in Table 16, our GRPO-based model tends to generate partial answers more than other models.
We also analyzed model performance on questions involving negation (e.g., those containing adverbs such as “not” or “never”). For questions that involve filtering-related skills, we have 132 instances with negation in our test set (Filtering: 11, Filtering+Aggregation: 15, Filtering+Sorting:12, and Relational Filtering:94). The results in Table 17 show that all models struggle with negation-based filtering.
We further manually analyzed 200 reasoning traces from reasoning models including Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates. We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis. For each SQL query template, we have different SQL queries and questions for different sizes of contexts, while the questions require the same reasoning capability. There are no reasoning traces when the input is 512K or 1M tokens for Qwen3-4B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B, as they do not support the input of 512K or 1M tokens.
Model A S F F+A F+S RF
Qwen2.5-7B-Instruct-1M 70 90 80 90 80 80
Qwen2.5-14B-Instruct-1M 10 10 20 50 50 60
Gemma-3-27B-IT 10 0 0 40 10 40
Llama-4-Scout-14Bx16E-Instruct 30 20 30 60 20 60
Llama-3.3-70B-Instruct 10 10 30 80 50 50
DeepSeek-R1-Distill-Llama-70B 0 0 0 10 0 10
Gemini 2.5 Pro 0 0 0 20 30 10
GPT-4.1 0 0 0 10 0 0
o4-mini 0 0 10 10 30 20
Qwen2.5-7B-Instruct-1M (SFT) 0 0 40 30 40 0
Qwen2.5-7B-Instruct-1M (GRPO) 0 0 0 0 0 0
Table 14: Proportion of samples (%) where models output “NULL” broken down across information processing skills. Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and RF for Relational Filtering.
Model A S F F+A F+S RF
Qwen2.5-7B-Instruct-1M 0 0 0 0 0 0
Qwen2.5-14B-Instruct-1M 0 30 10 0 30 0
Gemma-3-27B-IT 0 20 30 20 20 0
Llama-4-Scout-14Bx16E-Instruct 10 10 0 10 30 10
Llama-3.3-70B-Instruct 10 10 30 0 10 0
DeepSeek-R1-Distill-Llama-70B 0 0 0 0 0 0
Gemini 2.5 Pro 0 10 0 0 0 0
GPT-4.1 30 30 20 50 20 0
o4-mini 10 20 0 0 0 0
Qwen2.5-7B-Instruct-1M (SFT) 0 0 0 0 0 0
Qwen2.5-7B-Instruct-1M (GRPO) 0 0 0 10 0 0
Table 15: Proportion of samples (%) where models do not follow the specified answer format broken down across information processing skills. Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and RF for Relational Filtering.
Model A S F F+A F+S RF
Qwen2.5-7B-Instruct-1M 10 0 10 0 0 0
Qwen2.5-14B-Instruct-1M 0 0 0 0 0 10
Gemma-3-27B-IT 10 10 10 0 0 0
Llama-4-Scout-14Bx16E-Instruct 0 0 0 0 0 0
Llama-3.3-70B-Instruct 0 0 10 0 0 0
DeepSeek-R1-Distill-Llama-70B 0 0 0 20 0 0
Gemini 2.5 Pro 20 0 10 20 0 10
GPT-4.1 10 0 0 0 0 0
o4-mini 10 0 20 30 10 0
Qwen2.5-7B-Instruct-1M (SFT) 10 0 0 0 0 20
Qwen2.5-7B-Instruct-1M (GRPO) 20 0 30 10 20 0
Table 16: Proportion of samples (%) where models generate partial answers broken down across information processing skills. Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and RF for Relational Filtering.
Model F+A F+S F RF Avg
Qwen2.5-7B-Instruct-1M 0 0 0 0 0
Qwen2.5-14B-Instruct-1M 0 20.0 0 2.1 3.8
Gemma-3-27B-IT 0 6.7 0 0 0.8
Llama-4-Scout-14Bx16E-Instruct 0 20.0 0 2.1 3.8
Llama-3.3-70B-Instruct 0 6.7 0 0 0.8
DeepSeek-R1-Distill-Llama-70B 0 13.3 0 0 1.5
Gemini 2.5 Pro 0 20.0 8.3 6.4 9.1
GPT-4.1 0 13.3 0 2.1 3.8
o4-mini 0 73.3 41.7 7.4 22.0
Qwen2.5-7B-Instruct-1M (SFT) 0 6.7 0 11.7 9.1
Qwen2.5-7B-Instruct-1M (GRPO) 0 33.3 8.3 3.2 8.3
Table 17: Model performance on questions that involve negation (e.g., “not" or “never"), using exact match (%). Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and RF for Relational Filtering.
Appendix E Prompts used for SciTrek Creation and Model Evaluation
We generate natural language questions by prompting Qwen2.5-Coder-32B to convert SQL queries to natural language. Our prompt is presented in Figure 3. The prompt that we use to convert natural language questions back to to SQL queries is shown in Figure 4. We employ two different prompts for (zero-shot) model evaluation. The prompt for assessing LCLM capabilities against full-text articles is presented in Figure 5, and the prompt for using the corresponding database tables as context is shown in Figure 7. The prompt used by GRPO to generate reasoning races is given in Figure 6.
Instruction to Convert SQL Queries to Natural Language Questions You are given a database with three tables: articles, article-author, and citing-cited. The articles table contains the following columns: - article_\_id (String): the unique identifier of the article; - article_\_title (String): the title of the article; - title_\_word_\_count (Integer): the number of words in the article’s title (using spaces to determine word boundaries); - author_\_count (Integer): the number of authors for the article; - reference_\_count (Integer): the number of references cited in the article. The article-author table contains the following columns: - relation_\_id (String): the unique identifier of the article-author relationship; - article_\_id (String): the identifier of the associated article; - author_\_name (String): the name of the author; - author_\_position (Integer): the position of the author in the author list (starting from 0 for the first author). The citing-cited table contains the following columns: - relation_\_id (String): the unique identifier of the citation relationship between two articles; - article_\_id_\_citing (String): the identifier of the article which cites the other article; - article_\_id_\_cited (String): the identifier of the article which is cited by the other article. Assumptions: - The articles table contains multiple entries; - The article_author table maps authors to articles, where one author can contribute to multiple articles, and one article can have multiple authors; - The citing_cited table represents citation relationships among articles in the articles table, where one article can be cited by multiple others. Your task involves two steps: 1. Understand the given SQL query in the context of the database schema described above; 2. Convert the SQL query into a clear and natural-sounding question in everyday language, as if you were reading textural articles rather than querying a database. The given SQL query: {sql_query} Do not refer to relation_id or article_id in the natural-language question. You must output the SQL query and the corresponding question in the following JSON format, and do not include any extra text: {“sql": “the given SQL query", “question": “the generated question"}
Figure 3: Prompt template for converting SQL queries to natural language questions.
Instruction to Convert Natural Language Questions to SQL Queries You are given a database with three tables: articles, article-author, and citing-cited. The articles table contains the following columns: - article_\_id (String): the unique identifier of the article; - article_\_title (String): the title of the article; - title_\_word_\_count (Integer): the number of words in the article’s title (using spaces to determine word boundaries); - author_\_count (Integer): the number of authors for the article; - reference_\_count (Integer): the number of references cited in the article. The article-author table contains the following columns: - relation_\_id (String): the unique identifier of the article-author relationship; - article_\_id (String): the identifier of the associated article; - author_\_name (String): the name of the author; - author_\_position (Integer): the position of the author in the author list (starting from 0 for the first author). The citing-cited table contains the following columns: - relation_\_id (String): the unique identifier of the citation relationship between two articles; - article_\_id_\_citing (String): the identifier of the article which cites the other article; - article_\_id_\_cited (String): the identifier of the article which is cited by the other article. Assumptions: - The articles table contains multiple entries; - The article_author table maps authors to articles, where one author can contribute to multiple articles, and one article can have multiple authors; - The citing_cited table represents citation relationships among articles in the articles table, where one article can be cited by multiple others. Available core SQL commands: - Aggregating: MIN(), MAX(), COUNT(), SUM(), AVG(), DISTINCT - Filtering: WHERE - Organizing: ORDER BY, ASC, DESC, GROUP BY Available core SQL operators: - Comparison: ==, >>, <<, >=>=, <=<=, <><>, LIKE - Arithmetic: ++, -, *, //, % - Logical: AND, NOT, OR, BETWEEN, IN Your task is to: 1. Understand the database schema described above and the given natural language question below; 2. Convert the natural language question into a SQL query in the context of the database schema with the listed SQL commands and operators. The given natural language question: {question} Do not output relation_id or article_id in generated SQL query. Use the SQL commands and operators listed above. Make the generated SQL query aligned well with the natural language question. You must output the natural language question and the generated SQL query in the following JSON format, and do not include any extra text: {“question": “the given question", “sql": “the generated SQL query"}
Figure 4: Prompt template for converting natural language questions to SQL queries.
Testing Instruction Using Full-text Articles Articles: {scientific articles} You are provided with multiple scientific articles above. Based on the information in these articles, answer the question provided below. If the answer consists of multiple components (e.g., author names, article titles, reference counts), separate them with commas. For example, if the answer includes two author names, your response should be in the format of ‘the-first-author-name, the-second-author-name’. When counting the number of words in article titles, use spaces to determine word boundaries. Words are spaced apart individually. Respond with only the final answer, with no additional explanation or formatting. If you cannot get the answer from the articles, just return ‘NULL’. Question: {question}
Figure 5: Prompt template using full-text articles as context.
[…]
Figure 6: Prompt template to generate reasoning traces in reinforcement learning using full-text articles as context.
Testing Instruction Using Database Tables You are given three tables, named articles, article-author, and citing-cited. The articles table contains the following columns: - article_\_id (String): the unique identifier of the article; - article_\_title (String): the title of the article; - title_\_word_\_count (Integer): the number of words in the article’s title (using spaces to determine word boundaries); - author_\_count (Integer): the number of authors for the article; - reference_\_count (Integer): the number of references cited in the article. {the table of articles} The article-author table contains the following columns: - relation_\_id (String): the unique identifier of the article-author relationship; - article_\_id (String): the identifier of the associated article; - author_\_name (String): the name of the author; - author_\_position (Integer): the position of the author in the author list (starting from 0 for the first author). {the table of article-author} The citing-cited table contains the following columns: - relation_\_id (String): the unique identifier of the citation relationship between two articles; - article_\_id_\_citing (String): the identifier of the article which cites the other article; - article_\_id_\_cited (String): the identifier of the article which is cited by the other article. {the table of citing-cited} Based on the information in the tables above, answer the question provided below. If the answer consists of multiple components (e.g., author names, article titles, reference counts), separate them with commas. For example, if the answer includes two author names, your response should be in the format of ‘the-first-author-name, the-second-author-name’. Respond with only the final answer, with no additional explanation or formatting. If you cannot get the answer from the tables, just return ‘NULL’. Question: {question}
Figure 7: Prompt template using database tables as context.